# HiSTF Mamba: Hierarchical Spatiotemporal Fusion with Multi-Granular Body-Spatial Modeling for High-Fidelity Text-to-Motion Generation

### v1

@@ Abstract
Text-to-motion generation is a rapidly growing field at the nexus of multimodal learning and computer graphics, promising flexible and cost-effective applications in gaming, animation, robotics, and virtual reality. Existing approaches often rely on simple spatiotemporal stacking, which introduces feature redundancy, while subtle joint-level details remain overlooked from a spatial perspective. To this end, we propose a novel HiSTF Mamba framework. The framework is composed of three key modules: Dual-Spatial Mamba, Bi-Temporal Mamba, and Dynamic Spatiotemporal Fusion Module (DSFM). Dual-Spatial Mamba incorporates “Part-based + Whole-based” parallel modeling to represent both whole-body coordination and fine-grained joint dynamics. Bi-Temporal Mamba adopts a bidirectional scanning strategy, effectively encoding short-term motion details and long-term dependencies. DSFM further performs redundancy removal and extraction of complementary information for temporal features, then fuses them with spatial features, yielding an expressive spatio-temporal representation. Experimental results on the HumanML3D dataset demonstrate that HiSTF Mamba achieves state-of-the-art performance across multiple metrics. In particular, it reduces the FID score from 0.283 to 0.189, a relative decrease of nearly 30%. These findings validate the effectiveness of HiSTF Mamba in achieving high fidelity and strong semantic alignment in text-to-motion generation.
@@ 1 Introduction
To address these limitations in modeling fine-grained joint details and effectively fusing spatio-temporal features, we propose HiSTF Mamba. HiSTF Mamba consists of three key components: Dual-Spatial Mamba, Bi-Temporal Mamba, and DSFM. Concretely, Dual-Spatial Mamba adopts a parallel strategy “Part-based + Whole-based”, ensuring that the model can capture fine-grained joint details, while preserving coarse-grained whole-body coordination. Meanwhile, the Bi-Temporal Mamba block employs a bidirectional selected SSMs to scan the time dimension both forward and backward, capturing both short-term cues and long-term dependencies to ensure temporal coordination. Finally, rather than simply stacking spatiotemporal modules, we propose DSFM, which removes redundant temporal features from Bi-Temporal Mamba, extracts complementary information, and fuses it with the spatial output from Dual-Spatial Mamba for more efficient spatiotemporal fusion. Thus, HiSTF Mamba accommodates both global and local, long-range and short-range, as well as spatial and temporal dimensions, thereby significantly improving the fidelity and realism of text-to-motion generation.
We conducted extensive experiments on the HumanML3D [20] and KIT [46] datasets, achieving excellent results across multiple key evaluation metrics. The experimental findings indicate that the proposed HiSTF Mamba can generate motion sequences with a high degree of semantic alignment to the text, naturally smooth transitions, and enhanced visual realism. The results on the HumanML3D dataset are particularly remarkable: HiSTF Mamba demonstrates significant improvements in various evaluation metrics, reducing the FID from 0.281 to 0.189 compared to the current state-of-the-art method, an almost 30% decrease, while maintaining leading performance in other key metrics.
@@
    In the Dual-Spatial Mamba block, we propose a “Part-based + Whole-based” parallel strategy that captures both whole-body coordination and fine-grained joint-level details, thereby improving the model’s ability to represent complex human body structures.
@@
    Extensive experiments on HumanML3D demonstrate that HiSTF Mamba achieves outstanding results across multiple metrics. Notably, it reduces the FID score from 0.283 to 0.189, a reduction of nearly 30%, while also achieving leading performance on other metrics. This validates HiSTF Mamba’s superiority in balancing high fidelity and strong text relevance.
@@ 2.1 Text-to-Motion Generation
Early text-to-motion works could only generate limited motions, such as Text2Action [1], which is restricted to synthesizing upper-body motion sequences from text descriptions by combining Recurrent Neural Network (RNN) and GAN. Meanwhile, JL2P [2] achieves a constrained one-to-one mapping between text and motion by learning a joint embedding space for language and pose. To address the problem that a text can only correspond to a determined motion, Guo et al. [20] proposed a motion generation model based on conditional VAE, which can stochastically generate human motions of variable lengths. TEMOS [43] utilizes a transformer-VAE to learn a joint latent space for both motion and text modalities, achieving alignment between motion and text within this latent space. TEACH [4] further extends this capability to longer motion sequences. Similarly, MotionCLIP [50] directly aligns the motion space with the CLIP space [48], effectively integrating the rich semantic knowledge of CLIP into the motion representation. Subsequently, VQ-VAE [52] is introduced in TM2T [21] to discretize motion sequences into tokens. This discrete representation facilitates the subsequent text-to-motion translation process. These VQ-VAE-based methods [57, 45, 22, 44] have since dominated the text-to-motion task, significantly outperforming other types of methods on the evaluation metric FID. This success is largely attributed to the discrete token representation, where each token inherently represents a brief snippet of real motion, significantly reducing the discrepancy between generated motions and real motions. However, the trade-off is that representing continuous motion data as a finite set of discrete tokens inevitably leads to information loss, thereby reducing the diversity of the generated motions.
In addition to GAN and VAE, the diffusion model [24], as a generative model, has also been widely adopted in the field of motion generation. Diffusion models have demonstrated their ability to produce high-quality and diverse samples while ensuring stable training, as evidenced in the realm of text-to-image [49]. MotionDiffuse [59] is the first framework to apply diffusion models to text-driven motion synthesis. MDM [51] employs a transformer encoder as its denoiser, directly predicting clean samples rather than noise. Due to the time-consuming and inefficient nature of conducting diffusion natively in the original motion data domain, MLD [9] opts to perform the diffusion process within the latent space of motion. Recently, MotionLCM [13] has introduced consistency distillation [40] on the foundation of MLD to tackle the inefficiency of diffusion sampling, further accelerating motion generation to real-time levels. Unlike previous methods, EMDM [62] proposes sampling motion from the diffusion model with fewer sampling steps during the denoising phase, achieving significant acceleration without compromising motion quality. Inspired by Mamba [16], Motion Mamba [61] introduces the Bidirectional Spatial Mamba block and the Bidirectional Temporal Mamba block into the U-net architecture, aiming to enhance the consistency of motion generation.
@@ 2.2 State Space Models
Transformers [53] excel in sequence modeling but often incur prohibitive quadratic complexity for high-resolution vision tasks. In response, State Space Models (SSMs)[18, 19] perform linear recurrences to capture long-range dependencies. For example, LSSL[19] employs continuous-time linear memory, while S4 [18], leveraging HiPPO [17] and FFT-based convolutions, significantly reduces computational overhead. Building on S4, Mamba [16] introduces dynamic gating and a parallel scan, accelerating inference without sacrificing global context.
In vision tasks, SSM-based techniques alleviate Transformers’ high computational cost by serializing images or videos into manageable sequences [16]. For example, Vision Mamba [63] and VMamba [37] replace self-attention with scanning strategies to achieve performance comparable to Transformers, whereas PlainMamba [56] and LocalMamba [25] adopt continuous or local scans to diminish spatial discontinuities and refine fine-grained segmentation. EfficientVMamba [42] incorporates dilated convolutions and stride sampling, maintaining accuracy under stringent resource constraints. Beyond scanning alone, additional modules can bolster local feature extraction: convolutional branches [7, 23, 65] are integrated into MambaIR [23] to merge convolutional and SSM layers for improved image restoration. Furthermore, MMA [10] and VL-Mamba [47] interleave attention-based components with SSMs to excel in super-resolution and multimodal fusion, while SpikeMba [35] and VideoMamba [34] leverage SSMs for video generation and temporal modeling, efficiently handling lengthy sequences and reducing computational costs.
@@ 3 Method
Figure 1: (Left) HiSTF Mamba overview. The text condition, the noise step $t$, and the noise-augmented sequence $\{x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}\}$ are fed into the model. The text condition is encoded by the CLIP [48] text encoder, randomly masked, and projected into latent space to obtain the vector $c$. Next, $t$ is processed by an MLP and added to $c$ to form the token $Z_{\text{tk}}$. Finally, these inputs are used by the model to predict the denoised sequence $\{\hat{x}_{0}^{1},\hat{x}_{0}^{2},\hat{x}_{0}^{3},\ldots,\hat{x}_{0}^{N}\}$. (Right) Sample HiSTF Mamba. The text condition $c$, the noise step $t$, and a random noise $X_{T}$ are fed into the model to predict a clean sequence $\hat{x}_{0}$. The diffusion process then generates $x_{T-1}$ from $\hat{x}_{0}$. Repeating this procedure $T$ times yields a clean sequence $x_{0}$ that conforms to the text condition. DMP-Solver++ [39] is used to accelerate sampling.
In this section, we delineate the overall architecture and operational principles of the HiSTF Mamba, which aims to generate human motion guided by textual inputs. We begin by describing the theoretical foundations of Mamba [16], followed by an overview of the HiSTF Mamba. Subsequently, we detail the spatiotemporal modules: Dual-Spatial Mamba and Bi-Temporal Mamba, then proceed to discuss DSFM, which effectively fuses these spatiotemporal features. Finally, we introduce the loss function design for HiSTF Mamba.
@@ 3.1 Preliminaries
$\begin{cases}h_{t}=A\,h_{t-1}+B\,x_{t},\\[6.0pt] y_{t}=C\,h_{t},\end{cases}$ (1)
@@
$\begin{cases}h_{t}=A(\Delta_{t})\,h_{t-1}+B_{t}\,x_{t},\\[6.0pt] y_{t}=C_{t}\,h_{t},\end{cases}\quad\begin{cases}B_{t}&=s_{B}\bigl{(}x_{t}\bigr{)},\\ C_{t}&=s_{C}\bigl{(}x_{t}\bigr{)},\\ \Delta_{t}&=\tau_{\Delta}\!\bigl{(}W_{\Delta}\,x_{t}\bigr{)},\end{cases}$ (2)
@@ 3.2 HiSTF Mamba
Figure 2: (Left) Bi-Temporal Mamba. The input sequence is normalized, and then processed forward and backward to capture richer temporal dependencies for smooth, coherent motion. (Right) Dual-Spatial Mamba. The Part-based branch splits the full-body motion into six parts, extracts fine-grained local features via a 2D dynamic convolution and Part SSMs (Part Mamba), and recombines them in Part2Whole. Meanwhile, the Whole-based branch models the entire body (Whole SSMs) to learn globally coordinated patterns. Finally, local and global features are summed, obtaining a representation that is both well-coordinated and captures joint-level details.
Existing methods face two main issues. First, when the entire motion matrix is fed into the network at once, earlier joint features become overshadowed, leading to the loss of fine-grained details. Second, stacking spatio-temporal modules not only fails to effectively learn and fuse spatiotemporal features, resulting in duplicate or overlapping feature extraction that increases computational overhead and reduced fidelity. To address these problems, we propose HiSTF Mamba, as illustrated in Fig. 1 a hierarchical spatiotemporal fusion Mamba method for text-to-motion generation.
HiSTF Mamba addresses the difficulty of extracting fine-granularity joint details and the inability to effectively fuse spatio-temporal features through three specialized components: N Bi-Temporal Mamba blocks, one Dual-Spatial Mamba block, and a DSFM module. By modeling part-based and whole-based motions in parallel with Dual-Spatial Mamba and then dynamically fusing them with temporal features learned by Bi-Temporal Mamba using the DSFM module, HiSTF Mamba ensures whole-body coordination while preserving fine-grained joint details and efficiently fusing temporal features.
Our proposed HiSTF Mamba is built upon MDM [51]’s diffusion and sampling processes, and we integrate a DMP-Solver++ [39] strategy to accelerate sampling, reducing the original 1000 steps to 10 or 15. $c$, Specifically, HiSTF Mamba takes a processed text condition a noise step $t$, and a motion sequence $\{x_{t}^{1},x_{t}^{2},\dots,x_{t}^{N}\}$ subjected to $t$ diffusion steps, and outputs the corresponding clean sequence $\{x_{0}^{1},x_{0}^{2},\dots,x_{0}^{N}\}$. This design balances global motion patterns with fine-grained joint dynamics, producing more accurate reconstructions. Here, $x_{0}^{i}\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ is the dimensionality of each joint representation.
@@ 3.3 Bi-Temporal Mamba and Dual-Spatial Mamba
Figure 3: DSFM: Redundant information is removed and complementary information is extracted from the temporal features produced by the N Temporal Mamba blocks, yielding the refined temporal feature $\mathbf{T}$. This feature is then fused with the spatial feature $\mathbf{S}$ from the Space Mamba block to obtain a more expressive spatiotemporal feature representation – fused feature.
@@
Specifically, the input sequence is normalized. In the Part-based learning phase, we first map the normalized full-body motion sequence back to its original dimensionality. Drawing upon insights from ParCo [64], we design a Whole2Part module to split the whole-body motion into six parts. R.Leg, L.Leg, Backbone, Root, R.Arm, and L.Arm. This design follows the common practice of 3D human body models (e.g., SMPL, MMM), where Kinematic Trees typically represent the skeleton as five main chains (the four limbs plus the backbone); ParCo adds a separate Root component to account for global trajectories, thereby leading to six sub-motions in total. Formally, given a motion sequence $x=\{x^{1},x^{2},x^{3},\dots,x^{N}\}$, it is partitioned into $p_{i}=\{p_{i}^{1},p_{i}^{2},p_{i}^{3},\dots,p_{i}^{N}\},\quad i\in\{1,2,\ldots,S\}$, where $S$ denotes the total number of body parts. After segmenting the motion into six parts, the features of each part are projected to a latent dimension and fed into a 2D dynamic convolution network. To better capture the distinct motion patterns of each part, six learnable convolution kernels are used, each corresponding to a specific body part, thus improving the model’s capability to capture fine-grained local details. Next, we stack the outputs of these six parts along the batch dimension and feed them into the Selected SSMs (Part Mamba) module, which further models local dependencies and outputs refined local motion features. Finally, through the Part2Whole process, we restore the local characteristics to a holistic representation of the entire body, thus obtaining fine-grained characteristics at the joint level. In Whole-based learning phase, the full-body motion matrix is fed into a 1D convolutional network to extract key motion features of different parts. Subsequently, by leveraging the Selected SSMs (Whole SSMs) module’s advantages in sequence processing and long-range dependency modeling, inter-joint relationships and latent motion patterns at the global level are captured, thereby yielding global motion features. Finally, Part-based and Whole-based motion features are summed element-wise to obtain a representation that combines both global motion patterns and fine-grained joint-level detail features.

### v2

# Multi-granular body modeling with Redundancy-Free Spatiotemporal Fusion for Text-Driven Motion Generation

@@ Abstract
Text-to-motion generation is a rapidly growing field at the nexus of multimodal learning and computer graphics, promising flexible and cost-effective applications in gaming, animation, robotics, and virtual reality. Existing approaches often rely on simple spatiotemporal stacking, which introduces feature redundancy, while subtle joint-level details remain overlooked from a spatial perspective. To this end, we propose a novel HiSTF Mamba framework composed of three key modules: Dual-Spatial Mamba, Bi-Temporal Mamba, and the Dynamic Spatiotemporal Fusion Module (DSFM). Dual-Spatial Mamba incorporates “part-based + whole-based” parallel modeling to capture both whole-body coordination and fine-grained joint dynamics. Bi-Temporal Mamba employs a bidirectional scanning strategy that effectively encodes short-term motion details and long-term dependencies. DSFM further removes redundancy and extracts complementary information from temporal features, then fuses them with spatial features, yielding an expressive spatio-temporal representation. Experimental results on the HumanML3D dataset demonstrate that HiSTF Mamba achieves robust performance across multiple evaluation metrics; These findings validate the effectiveness of HiSTF Mamba in achieving high fidelity and strong semantic alignment in text-to-motion generation.
@@ 1 Introduction
Figure 1: Qualitative visualization of HiSTF Mamba’s generated motions.
@@
To address these limitations in modeling fine-grained joint details and effectively fusing spatio-temporal features, we propose HiSTF Mamba. HiSTF Mamba consists of three key components: Dual-Spatial Mamba, Bi-Temporal Mamba, and DSFM. Concretely, Dual-Spatial Mamba adopts a parallel strategy “Part-based + Whole-based”, ensuring that the model can capture fine-grained joint details, while preserving coarse-grained whole-body coordination. Meanwhile, the Bi-Temporal Mamba block employs a bidirectional selected SSMs to scan the time dimension both forward and backward, capturing both short-term cues and long-term dependencies to ensure temporal coordination. Finally, rather than simply stacking spatiotemporal modules, we propose DSFM, which removes redundant temporal features from Bi-Temporal Mamba, extracts complementary information, and fuses it with the spatial output from Dual-Spatial Mamba for more efficient spatiotemporal fusion. Thus, HiSTF Mamba accommodates both global and local, long-range and short-range, as well as spatial and temporal dimensions, thereby significantly enhancing the fidelity and realism of text-to-motion generation. As evidenced by the visualizations in Fig 1, the synthesized motions not only closely align with the text descriptions but also exhibit full-body coordination and joint-level detail, reflecting the model’s effectiveness in capturing nuanced spatiotemporal patterns.
@@
    In the Dual-Spatial Mamba block, we propose a “Part-based + Whole-based" parallel strategy that captures both whole-body coordination and fine-grained joint-level details, thereby improving the model’s ability to represent complex human body structures.
@@
    Extensive experiments on HumanML3D demonstrate that HiSTF Mamba achieves outstanding results across multiple metrics. Notably, it reduces the FID score from 0.281 to 0.189, a reduction of nearly 30%, while also achieving leading performance on other metrics. This validates HiSTF Mamba’s superiority in balancing high fidelity and strong text relevance.
@@ 2.1 Text-to-Motion Generation
Early text-to-motion methods offered limited diversity: Text2Action [26] generated only upper-body motion, while JL2P [27] mapped text to poses one-to-one. Guo et al. [28] introduced a conditional VAE for variable-length outputs. TEMOS [29] employed a transformer-VAE for unified motion-text representations, and TEACH [30] extended this idea to longer sequences. MotionCLIP [31] aligned motion with CLIP [32], while TM2T [33] used VQ-VAE [34] to discretize motion into tokens. Although VQ-VAE-based methods [35, 36, 37, 38] excel on FID, they risk information loss from quantization. Beyond GANs and VAEs, diffusion models [17] have proven adept at producing high-quality, diverse motions, exemplified by MotionDiffuse [20] and MDM [18]. MLD [19] moved diffusion to a latent space, and MotionLCM [39] added consistency distillation [40] for real-time generation. EMDM [41] further speeds sampling by reducing diffusion steps without degrading motion quality.
@@ 2.2 State Space Models
Transformers [42] excel in sequence modeling but often incur prohibitive quadratic complexity for high-resolution vision tasks. In response, State Space Models (SSMs)[43, 44] perform linear recurrences to capture long-range dependencies. For example, LSSL[44] employs continuous-time linear memory, while S4 [43], leveraging HiPPO [45] and FFT-based convolutions, significantly reduces computational overhead. Building on S4, Mamba [25] introduces dynamic gating and a parallel scan, accelerating inference without sacrificing global context. In vision tasks, SSM-based methods reduce Transformers’ computational cost by serializing images or videos into manageable sequences [25]. Vision Mamba [46] and VMamba [47] replace self-attention with scanning strategies, achieving Transformer-level performance. PlainMamba [48] and LocalMamba [49] use continuous or local scans to mitigate spatial discontinuities and improve segmentation. EfficientVMamba [50] further enhances efficiency with dilated convolutions and stride sampling, preserving accuracy under limited resources. Beyond scanning alone, additional modules can bolster local feature extraction: convolutional branches [51, 52, 53] are integrated into MambaIR [52] to merge convolutional and SSM layers for improved image restoration. Furthermore, SpikeMba [54] and VideoMamba [55] leverage SSMs for video generation and temporal modeling, efficiently handling lengthy sequences and reducing computational costs.
@@ 3 Method
In this section, we delineate the overall architecture and operational principles of the HiSTF Mamba, which aims to generate human motion guided by textual inputs. We begin by describing the theoretical foundations of Mamba [25], followed by an overview of the HiSTF Mamba. Finally, we detail the spatiotemporal modules: Dual-Spatial Mamba and Bi-Temporal Mamba, then proceed to discuss DSFM, which effectively fuses these spatiotemporal features.
@@ 3.1 Preliminaries
$\begin{cases}h_{t}=A\,h_{t-1}+B\,x_{t},\\[6.0pt] y_{t}=C\,h_{t}.\end{cases}$ (1)
@@
$\begin{cases}h_{t}=A(\Delta_{t})\,h_{t-1}+B_{t}\,x_{t},\\[6.0pt] y_{t}=C_{t}\,h_{t},\end{cases}\quad\begin{cases}B_{t}&=s_{B}\bigl{(}x_{t}\bigr{)},\\ C_{t}&=s_{C}\bigl{(}x_{t}\bigr{)},\\ \Delta_{t}&=\tau_{\Delta}\!\bigl{(}W_{\Delta}\,x_{t}\bigr{)}.\end{cases}$ (2)
@@ 3.2 HiSTF Mamba
To address the lack of local body structure modeling and suboptimal spatiotemporal fusion in text-to-motion generation, we propose HiSTF Mamba, a hierarchical spatiotemporal fusion approach, as illustrated in Fig. 2. Through three specialized components: N Bi-Temporal Mamba blocks, a Dual-Spatial Mamba block, and a DSFM module, HiSTF Mamba effectively captures fine-grained joint details and fuses spatiotemporal features. Specifically, the input sequence is first processed by N layers of Bi-Temporal Mamba to model rich temporal dependencies, and the resulting features are then fed into the Dual-Spatial Mamba block. Within this block, 2D dynamic convolution and Part SSMs extract joint-detail features from six body segments, which are then combined with the Whole SSMs’ global coordination to obtain the spatial feature $S$. Next, the N temporal features $\{T_{1},\ldots,T_{N}\}$ are passed to the DSFM for redundancy removal and subsequently fused with $S$, yielding a spatiotemporal motion representation that preserves both joint-level detail and overall body coordination.
In the sampling stage, we first encode the text condition with the CLIP [32] text encoder, apply random masking, and project it into the latent space. Meanwhile, we project the noise step $T$ through an MLP and add it to the encoded text condition, obtaining $x^{(N+1)}$. we incorporate the $x^{(N+1)}$ with the motion sequence subjected to $T$ steps of noise, yielding $\{x_{T}^{1},x_{T}^{2},\ldots,x_{T}^{N}\}$, which is fed into the model to predict the clean sequence $\{\hat{{x}}_{0}^{1},\hat{{x}}_{0}^{2},\dots,\hat{{x}}_{0}^{N}\}$. The diffusion process then generates $x^{T-1}$ from $\{\hat{{x}}_{0}^{1},\hat{{x}}_{0}^{2},\dots,\hat{{x}}_{0}^{N}\}$. Repeating this procedure $T$ times ultimately produces a clean sequence $\{{{x}}_{0}^{1},{{x}}_{0}^{2},\dots,{{x}}_{0}^{N}\}$ that conforms to the text condition. This design balances global motion patterns with fine-grained joint dynamics, producing more accurate reconstructions. DPM-Solver++ [56] is used to accelerate sampling.
Figure 2: The overview of the proposed HiSTF Mamba. (a) Sampling Process. The model receives a text prompt as input and begins the iterative generation process from Gaussian noise, ultimately producing a noise-free action sequence $x_{0}$ that is semantically aligned with the prompt. (b) HiSTF Mamba Module. The HiSTF Mamba module primarily aims to predict a clean motion sequence $\{\hat{{x}}_{0}^{1},\hat{{x}}_{0}^{2},\dots,\hat{{x}}_{0}^{N}\}$ from a noisy input sequence $\{x_{T}^{1},x_{T}^{2},\ldots,x_{T}^{N}\}$. By leveraging Bi-Temporal Mamba and Dual-Spatial Mamba, the model extracts temporal features $\{T_{1},\dots,T_{N}\}$ and a spatial feature $S$. These $N$ temporal features are then fed into the DSFM module to remove redundancy and are subsequently fused with $S$, producing a spatiotemporal motion representation that captures both joint-level details and overall body coordination. About Whole2Parts and Parts2Whole, see the appendix A.4
@@ 3.3 Bi-Temporal Mamba and Dual-Spatial Mamba
Specifically, the input sequence is normalized. In the Part-based learning phase, we first map the normalized full-body motion sequence back to its original dimensionality. Drawing upon insights from ParCo [57], we design a Whole2Part module to split the whole-body motion into six parts. R.Leg, L.Leg, Backbone, Root, R.Arm, and L.Arm. This design follows the common practice of 3D human body models (e.g., SMPL, MMM), where Kinematic Trees typically represent the skeleton as five main chains (the four limbs plus the backbone); We adds a separate Root component to account for global trajectories, thereby leading to six sub-motions in total. Formally, given a motion sequence $x=\{x^{1},x^{2},x^{3},\dots,x^{N}\}$, it is partitioned into $p_{i}=\{p_{i}^{1},p_{i}^{2},p_{i}^{3},\dots,p_{i}^{N}\},\quad i\in\{1,2,\ldots,S\}$, where $S$ denotes the total number of body parts. For more details about Whole2Parts and Parts2Whole, see the appendix A.4. After segmenting the motion into six parts, the features of each part are projected to a latent dimension and fed into a 2D dynamic convolution network. To better capture the distinct motion patterns of each part, six learnable convolution kernels are used, each corresponding to a specific body part, thus improving the model’s capability to capture fine-grained local details. Next, we stack the outputs of these six parts along the batch dimension and feed them into the Selected SSMs (Part Mamba) module, which further models local dependencies and outputs refined local motion features. Finally, through the Part2Whole process, we restore the local characteristics to a holistic representation of the entire body, thus obtaining fine-grained characteristics at the joint level. In Whole-based learning phase, the full-body motion matrix is fed into a 1D convolutional network to extract key motion features of different parts. Subsequently, by leveraging the Selected SSMs (Whole SSMs) module’s advantages in sequence processing and long-range dependency modeling, inter-joint relationships and latent motion patterns at the global level are captured, thereby yielding global motion features. Finally, Part-based and Whole-based motion features are summed element-wise to obtain a representation that combines both global motion patterns and fine-grained joint-level detail features.